Monday, March 6, 2000

Mathematical Issues in the Sciences: An Exploratory Workshop

Center for Nonlinear Analysis Carnegie Mellon University Pittsburgh

Scope: This workshop will draw together researchers from CMU working on diverse research areas in Sciences and Engineering, all using sophisticated mathematical tools and techniques in their research programs. Speakers will be asked to set up the problems in their physical context, and to identify the main mathematical challenges that they are facing. There will be ample time for discussion.

Speakers:

Marc DeGraef -  Department of Materials Science & Engineering
Martial Hebert -  Robotics Institute
Myung Jhon -  Department of Chemical Engineering
Ramamoorthi Ravi - Graduate School of Industrial Administration
Larry Wasserman - Department of Statistics
Michael Widom -  Department of Physics

Location: Baker Hall, Adamson Wing

If you are planning to attend, please register by contacting Nancy Watson at nw0z@andrew.cmu.edu or x82545 ON OR BEFORE Tuesday, February 29.

Organizers: Irene Fonseca, Reha Tütüncü, Noel Walkington

For further information please contact: Mathematical Issues in the Sciences: An Exploratory Workshop Center for Nonlinear Analysis Carnegie Mellon University Department of Mathematical Sciences Pittsburgh, PA 15213 Telephone: (412) 268-2545 Fax:

(412) 268-6380 E-mail: cn0s@andrew.cmu.edu http://www.math.cmu.edu/cna

The CNA, established in 1991, is a National Science Foundation sponsored research and training center concentrating in areas including nonlinear analysis, mechanics, scientific computation, and mathematical finance.